Find an equation of the line parallel to 5x + 4y = 2 containing the point (3, –1).
Answer by lwsshak3(11628)

Find an equation of the line parallel to 5x + 4y = 2 containing the point (3, –1)
**
Equation of a line: y=mx+b, m=slope, b=y-intercept
5x + 4y = 2
4y=-5x+2
y=-5x/4+1/2
m=-5/4
..
lines parallel to each other have the same slopes
equation of parallel line:
y=-5x/4+b
solve for b using coordinates of given point (3,-1)
-1=-5*3/4+b
b=-1+15/4=-4/4+15/4=11/4
equation: y=-5x/4+11/4